# People’s Perceptions Toward Bias and Related Concepts in Large Language Models: A Systematic Review

### v1

@@ 5.3. Implications for CHI
Despite the potential, successful adoption of these tools would require people’s trust in these systems. However, our review found people’s concerns toward LLM and the various aspects of its performance, including coherence, appropriateness, and explainability. Given that LLMs are new and still continuously evolving, continued research should be done to further understand how people’s perceptions of these systems also evolve together, and how these negative perceptions can be addressed through novel design solutions. For instance, prompt engineering requires a complex understanding of the model, and users met challenges in generating, evaluating, and explaining prompts’ effects (Zamfirescu-Pereira et al., 2023), indicating the necessity of understanding users’ mental models of how LLMs work and devising solutions to address user perceptions that might generate challenges around over- or under-trust of these systems.

### v2

@@ 5.3. Implications for CHI
Despite the potential, successful adoption of these tools would require people’s trust in these systems. Researchers define human-centered AI as supporting its human users “while revealing its underlying values, biases, limitations, and the ethics of its data gathering and algorithms to foster ethical, interactive, and contestable use” (Capel and Brereton, 2023), we are still at an early stage of testing how people perceive and experience bias in LLMs and being able to communicate these limitations and biases to users effectively (Xu et al., 2019; Liao et al., 2020). Our review found people’s concerns toward LLM and the various aspects of its performance, including coherence, appropriateness, and explainability. Given that LLMs are new and still continuously evolving, continued research should be done to further understand how people’s perceptions of these systems also evolve together, and how these negative perceptions can be addressed through novel design solutions. For instance, prompt engineering requires a complex understanding of the model, and users met challenges in generating, evaluating, and explaining prompts’ effects (Zamfirescu-Pereira et al., 2023), indicating the necessity of understanding users’ mental models of how LLMs work and devising solutions to address user perceptions that might generate challenges around over- or under-trust of these systems.